# Partially stochastic deep learning with uncertainty quantification for model predictive heating control

### v2

@@ 3.2 Model implementations
Given the unique thermodynamics of each building and its heating system, models are trained separately for each of the $S$ buildings. Thus, we end up having $S$ reference models, $S$ LSTM+MLP models, and $S$ LSTM+BNN models. All model formulations can be extended to any building with sufficient data on inputs and outputs, allowing the building-specific thermal dynamics to be expressed through model parameters learned from the data. The DL models are flexible enough to train a single global model for all buildings, enabling the model to transfer information between buildings with similar thermal behavior [9]. While this approach may benefit buildings with insufficient datasets, our aim in this study is to develop building-specific models for cases with comprehensive datasets. In industrial heating MPC applications, one can continuously evaluate different models and the available datasets for different buildings and prediction circumstances, and dynamically choose the model with the best potential for each building and prediction scenario.
Details regarding the inputs and outputs of the reference model are covered in Section 2.2. The state and the parameter estimation of the reference model for each building is performed using the data of the respective building from the last 500 heating season days prior to September 2021.
To train the DL models, all available site-specific data prior to September 2021 are split into training and validation sets at a 9-to-1 ratio. The models are implemented using software packages PyTorch [43] and torchbnn [44] as auto-regressive predictors that work on an hourly domain. That is, although the models are used to predict multiple time steps into the future, in the training phase, they are fitted based on their 1-hour forward predictions [9]. Both DL model variants are trained using the Adam optimizer with an initial learning rate of $10^{-4}$, decayed by a factor of $\frac{1}{2}$ over 3 scheduled reductions. The number of training epochs are 400 for the LSTM+MLP and 800 for the LSTM+BNN.
@@ 3.3 Results
The $\Delta T_{{in}}$ predictions of the DL models are transformed into $T_{{in}}$ predictions by cumulatively adding the predicted differences to the latest indoor temperature measurement the model had available given a starting point $t\in\bm{t}$. Since $T_{{in}}$ is incorporated in two of the DL models’ input variables (see Table 1), after the first prediction, the other predictions into the prediction horizon are affected by the previous predictions. In the reference model, the same applies inherently due to the LSSM formulation.
@@ 3.3.1 Prior selection for the partially stochastic deep learning model
The respective RMSE distributions are shown in Figure 3.
Figure 3: The performance of the LSTM+BNN with different prior variances $\beta^{2}$ as RMSE (17) for different prediction horizon lengths. Orange lines are the RMSE distribution medians, while boxes and whiskers indicate the $50\%$ and $95\%$ confidence intervals, respectively. Note the different scales on the Y-axis.
In addition to predictive performance, the value at which the KL divergence (2) converges during training must be considered. Since KL divergence measures the dissimilarity between the true and the approximate distribution, the closer we get to zero, the better the posterior approximation. Although the LSTM+BNN with a prior variance of $10^{-4}$ has the best accuracy, its KL divergence for the posterior approximation is 100 times larger than with prior variances of $10^{-2}$ and $10^{-3}$. A small variance in combination with a large KL divergence suggests that the prior is too restrictive. This can lead to instability or overfitting, as the model may not effectively regularize its parameters. Therefore, since the prior $\mathcal{N}(0,10^{-3}\bm{I})$ achieves good predictive performance and low KL divergence for the posterior approximation, it is selected for the LSTM+BNN employed in the comparative experiments of the study.
@@ 3.3.2 Predictive performance
For two test samples, Figure 4 shows the predictions of all three models and the predictive uncertainties of the LSTM+BNN model over the 48-hour prediction horizon. Since the uncertainty of the successive LSTM+BNN model predictions are independent, the predictive uncertainties over the prediction horizon are calculated by accumulating the standard deviations of preceding predictions in the sequence.
Figure 4: a) and b): 48-hour indoor temperature predictions for two different test samples. c) and d): Predictive uncertainties of the LSTM+BNN for the respective test samples on the left. Uncertainty is quantified as the cumulative standard deviation across each predictions’ 10 posterior predictive samples.
To compare the predictive performance of the models across different buildings and heating season conditions, we use the same RMSE metric (17) as we did when selecting the prior variance for the LSTM+BNN in 3.3.1. The results are shown in Figure 5. Both DL models outperform the reference for all three prediction horizon lengths, particularly for longer prediction horizons. The short-term predictions of the DL models are also significantly more accurate with a tighter RMSE distribution. The predictive performance between the LSTM+MLP and LSTM+BNN models is approximately equal.
Figure 5: Performance of the three models as RMSE (17) distributions into different prediction horizon lengths. Orange lines are the RMSE distribution medians, while boxes and whiskers indicate the $50\%$ and $95\%$ confidence intervals, respectively. Note the different scales on the Y-axis.
@@ 3.3.3 Uncertainty quantification
Figure 8 illustrates the average increase in prediction error of the LSTM+BNN with respect to predictive uncertainty. Although the LSTM+BNN performs slightly worse in terms of pure predictive performance compared to its deterministic variant, LSTM+MLP, (see Figure 6 and Table 2), it offers a key advantage in having the ability to quantify the uncertainty present in its predictions. This feature of offering transparency about the potential of the model is highly important when it comes to dynamically selecting the best model for each building and prediction scenario in industrial heating MPC applications.
Figure 8: Prediction uncertainties of the LSTM+BNN as standard deviations and the respective prediction errors as MAE for the $S\cdot T$ predictions into a 1-hour prediction horizon. The predictions and the standard deviations for a single test sample are calculated by computing a total of 100100 predictions using the building-specific LSTM+BNN models, and taking the mean and the standard deviation, respectively.
@@ 3.3.4 Computational costs
The primary drawback of DL models, compared to the reference model, is their computational cost (see Table 3). In the industrial heating MPC application for which the models are developed, control optimization—where models are used for predictions—and model training are separate tasks. For each building, model training is scheduled every other hour, while control optimization is performed hourly.
@@
Table 3: Computational costs of the models. The training time is the average time taken to train a model for a single building. The prediction time is the average time it took to generate a single 48-hour prediction sequence.
Given that the reference model is trained using data from the last 500 heating season days, such frequent model updates are not necessary to accurately reflect the recent thermodynamics of a building. Nonetheless, these updates are currently conducted so often simply due to the minimal computational cost of the reference model’s posterior inference. When integrating the DL models into the MPC framework, it would be adequate to train the model, for instance, only once a week, for buildings with sufficiently comprehensive datasets.
@@
During control optimization, the space heating supply temperature control sequence for the next $H$ hours is initialized based on the most recent optimization results. This sequence is optimized according to various objectives, taking into account recent observations of the system state, current weather forecasts, and model predictions. Due to the effective initialization, the optimization process typically concludes after just a few iterations. Consequently, the longer prediction time of the LSTM+BNN, compared to the reference model, does not impede the implementation of these models in control optimization.
4 Conclusions
We developed a partially stochastic DL architecture, termed LSTM+BNN, for building-specific indoor temperature modeling that incorporates built-in predictive uncertainty quantification.
The method was compared against its deterministic DL variant, LSTM+MLP, and a physics-based Bayesian LSSM used widely in a industrial MPC solution. Experiments were conducted using a large dataset collected from 100 centrally-heated residential buildings, covering different heating season conditions. Both of the DL models outperformed the reference model significantly in both short- and long-term predictions. The 48-hour predictions of the LSTM+BNN models were, on average, 42–43 % more accurate than those of the reference models, depending on the weighting of the short- versus long-term predictions. The improved performance of the DL models comes from their ability to capture complex nonlinear thermodynamics, including thermal lag effects (heat storage in building structures) and solar energy gain variations.
@@ 4 Conclusions
While the two DL variants showed similar accuracy, the partially stochastic version offers transparency by providing a way of quantifying the uncertainty present in its predictions. This feature is highly beneficial for integrating the data-driven model into real-world heating MPC solutions, as model suitability for a building and prediction scenario can be dynamically evaluated prior to each control optimization.
@@
The computational cost of the LSTM+BNN is much higher compared to that of the reference method. However, this does not pose a problem since the models could be trained less frequently when the dataset is comprehensive enough. Therefore, integrating the partially stochastic LSTM+BNN into heating MPC solutions — especially for buildings with comprehensive datasets — has great potential to improve control accuracy, thereby enhancing heating control in terms of energy efficiency, thermal comfort, and more advanced optimization objectives.
@@
Finally, the results of this work motivate and lay the groundwork for further studies on improving indoor temperature modeling related to heating MPC. Hence, we leave as future work the following:
@@
    Understanding carefully which of the learned physical relationships from the data contribute to the enhanced modeling ability of data-driven models, i.e., identifying what is missing from the existing reference model formulation and how it could be improved. For instance, methods like [45] could be used to find interpretable, functional formulations for the input-output mappings learned by the data-driven models.
@@
    Further optimization of the partially stochastic LSTM+BNN architecture (e.g., dimension reduction via sensitivity analysis or principal component analysis), and training settings via hyperparameter tuning (batch size, learning rates, initialization) to lower the computational costs.

### v3

@@ 3.2 Model implementations
Given the unique thermodynamics of each building and its heating system, models are trained separately for each of the $S$ buildings. Thus, we end up having $S$ reference models, $S$ LSTM+MLP models, and $S$ LSTM+BNN models. All model formulations can be extended to any building with sufficient data on inputs and outputs, allowing the building-specific thermal dynamics to be expressed through model parameters learned from the data. Although the DL models are flexible enough to train a single global model for knowledge transfer between buildings with similar thermal behavior [9]—an approach beneficial for buildings with insufficient data—our aim here is to develop building-specific models for cases with comprehensive datasets. In industrial heating MPC applications, one can continuously evaluate different models and the available datasets for different buildings and prediction circumstances, and dynamically choose the model with the best potential for each building and prediction scenario.
Details regarding the inputs and outputs of the reference model are covered in Section 2.2. The state and parameter estimation of the reference model for each building is performed using the data of the respective building from the last 500 heating season days prior to September 2021.
To train the DL models, all available site-specific data prior to September 2021 are split into training and validation sets at a 9-to-1 ratio. The models are implemented using software packages PyTorch [53] and torchbnn [54] as auto-regressive predictors that work on an hourly domain. That is, although the models are used to predict multiple time steps into the future, in the training phase, they are fitted based on their 1-hour forward predictions [9]. Both DL model variants are trained using the Adam optimizer with an initial learning rate of $10^{-4}$, decayed by a factor of $\frac{1}{2}$ over 3 scheduled reductions. The number of training epochs is 400 for the LSTM+MLP and 800 for the LSTM+BNN. Both models are trained using a single batch, and no specific random seed is specified.
@@ 3.3 Results
The $\Delta T_{{in}}$ predictions of the DL models are transformed into $T_{{in}}$ predictions by cumulatively adding the predicted differences to the latest indoor temperature measurement the model had available given a starting point $t\in\bm{t}$. Since $T_{{in}}$ is incorporated in the DL models’ input variables (see Table 1), each prediction beyond the first one into the prediction horizon is affected by the model’s own previous predictions. In the reference model, the same applies inherently due to the LSSM formulation.
Finally, we note that the predictions of the probabilistic LSSM and LSTM+BNN models refer to their respective predictive means. The Monte Carlo sampling strategy for generating probabilistic predictions from the partially stochastic LSTM+BNN is detailed in Section 2.3.2. In all numerical experiments, we used $N_{S}=10$ posterior predictive samples, with the exception of Section 3.3.3, for which $N_{S}=100$.
@@ 3.3.1 Prior selection for the partially stochastic deep learning model
The respective RMSE distributions are shown in Figure 3. Notice the different scales on the Y-axes to facilitate easier visual comparison of the LSTM+BNN with different prior variances for each prediction horizon length $K$.
Figure 3: Performance of the LSTM+BNN with different prior variances $\beta^{2}$ as RMSE (17) for different prediction horizon lengths $K\in[1,6,48]$. Orange lines are the RMSE distribution medians, while boxes and whiskers indicate the central $50\%$ and $95\%$ of the data, respectively. Note the different scales on the Y-axes.
In addition to predictive performance, the value at which the KL divergence term in (16) converges during training must be considered. Since this term measures the dissimilarity between the prior and the variational approximation, the closer we get to zero, the better. Although the LSTM+BNN with a prior variance of $10^{-4}$ has the best accuracy, the KL divergence term is 100 times larger compared to the models with prior variances of $10^{-2}$ and $10^{-3}$. A small prior variance combined with a large KL divergence suggests that the prior is too restrictive. This can lead to instability or overfitting, as the model may not effectively regularize its parameters. Therefore, since the prior $\mathcal{N}(0,10^{-3}\bm{I})$ achieves both good predictive performance and a low value for the KL divergence term, it is selected for the LSTM+BNN employed in the comparative experiments of the study.
This empirical analysis provides a principled, data-driven justification for the selected prior variance of $10^{-3}$, balancing predictive accuracy with stable posterior regularization for the building types in our dataset. We note that the optimal value is inherently dependent on the data characteristics of the specific buildings under study. A comprehensive sensitivity analysis across a wider variety of building types and a deeper physical interpretation of these hyperparameters would be greatly facilitated by a hierarchical Bayesian framework, where the priors themselves are learned from the data. This is a promising direction for enhancing the model’s generalizability and robustness, which we identify as an important objective for future work.
Finally, Table 2 summarizes all hyperparameters of the DL models employed in the further experiments of this study.
Table 2: Hyperparameters of the deep learning models.
LSTM+MLP LSTM+BNN
Epochs 400 800
Initial learning rate $10^{-4}$
Optimizer Adam
Hidden units 1024 / 1024 / 512
KL weight, $w_{\rm KL}$ - $10^{-3}$
Prior variance, $\beta^{2}$ - $10^{-3}$
@@ 3.3.2 Predictive performance
For two test samples, Figure 4 shows the predictions of all three models and the predictive uncertainties of the LSTM+BNN model over the 48-hour prediction horizon. Since the uncertainties of the successive LSTM+BNN model predictions are independent, the predictive uncertainty at each prediction horizon length $K\in[1,\ldots,48]$ is calculated by accumulating the standard deviations of all predictions up to that point.
Figure 4: a) and b): 48-hour indoor temperature predictions for two different test samples. c) and d): Predictive uncertainties of the LSTM+BNN corresponding to the test samples shown in a) and b), respectively. Uncertainty is quantified as the cumulative standard deviation across each prediction’s $N_{S}=10$ posterior predictive samples.
To compare the predictive performance of the models across different buildings and heating season conditions, we use the same RMSE metric (17) as we did when selecting the prior variance for the LSTM+BNN in Section 3.3.1. The results are shown in Figure 5. Again, notice the different scales on the Y-axes to facilitate easier visual comparison of the models for each prediction horizon length $K\in[1,6,48]$. Both DL models outperform the reference for all three prediction horizon lengths, particularly for longer prediction horizons. The short-term predictions of the DL models are also significantly more accurate with a tighter RMSE distribution. The difference in predictive performance between the LSTM+MLP and LSTM+BNN models is minimal.
Figure 5: Performance of the three models as RMSE (17) distributions into different prediction horizon lengths $K\in[1,6,48]$. Orange lines are the RMSE distribution medians, while boxes and whiskers indicate the central $50\%$ and $95\%$ of the data, respectively. Note the different scales on the Y-axes.
@@ 3.3.3 Uncertainty quantification
Figure 8 illustrates the average increase in prediction error of the LSTM+BNN models with respect to the predictive uncertainty. Although the LSTM+BNN performs slightly worse in terms of pure predictive performance compared to its deterministic variant, LSTM+MLP, (see Figure 6 and Table 3), it offers a key advantage in having the ability to quantify the uncertainty present in its predictions. This transparency regarding model reliability is highly beneficial when it comes to dynamically selecting the best model for each building and prediction scenario in industrial heating MPC applications.
Figure 8: Predictive uncertainties of the LSTM+BNN models (as standard deviations) and errors of the respective predictive means (as MAEs) for the $S\cdot T$ test samples into a 1-hour prediction horizon. The predictive mean and standard deviation were calculated from $N_{S}=100$ posterior predictive samples for each test sample.
@@ 3.3.4 Computational costs
The primary consideration for deploying DL models, compared to the reference model, is their computational cost (see Table 4). In the industrial heating MPC application for which the models are developed, the workflows for model training, model selection, and control optimization—where models are used for predictions—are separate and parallelizable computational tasks. This architecture is key to scalability. Currently, for each building, the reference model training is scheduled for every other hour, the model selection for once a week, and the control optimization is performed hourly.
Table 4: Computational costs of the models. The training time is the average time taken to train a model for a single building from scratch. The prediction time is the average time it took to generate a single 48-hour prediction sequence.
@@
For buildings with limited data, such frequent model updates are typically necessary, making the reference model with computationally efficient posterior inference the preferred choice. However, for buildings with comprehensive datasets where the DL models are applicable, the training frequency can be significantly reduced—for instance, to once per week. Furthermore, the reported training times (Table 4) are for training the models from scratch. In practice, retraining can be initialized from the previous model’s parameters, drastically reducing the number of iterations and computational cost required for convergence. The proposed LSTM+BNN model is intended for MPC of these data-rich buildings, where its superior accuracy justifies the computational expense. Therefore, while the aggregate training cost across a large portfolio is a defined limitation, the appropriate retraining scheduling and initialization enabled by the scalable architecture prevent it from being a prohibitive barrier to industrial deployment.
@@
In the control optimization task, the space heating supply temperature control sequence for the next $H$ hours is initialized based on the most recent optimization results. This sequence is then optimized according to various objectives, taking into account recent observations of the system state, current weather forecasts, and model predictions. Due to the effective initialization, the optimization process typically concludes after just a few iterations. Consequently, the slightly longer prediction time of the LSTM+BNN compared to the reference model does not hinder its practical use in the control optimization loop.
4 Conclusions and future work
We developed a partially stochastic DL architecture, termed LSTM+BNN, for building-specific indoor temperature modeling that incorporates built-in UQ. The method was compared against its deterministic DL variant, LSTM+MLP, and an industry-proven physics-based Bayesian LSSM. Rigorous comparative experiments were conducted using a large dataset collected from 100 centrally-heated residential buildings, extensively covering different heating season conditions.
@@ 4 Conclusions and future work
Both of the DL models outperformed the reference model significantly in short- and long-term predictions. The improved performance of the DL models comes from their ability to capture complex nonlinear thermodynamics, including thermal lag effects (heat storage in building structures) and solar energy gain variations. While the DL variants showed similar accuracy, the partially stochastic version offers transparency by providing a way of quantifying the uncertainty present in its predictions. This feature is highly beneficial for integrating the data-driven model into real-world heating MPC solutions, as model suitability for a building and prediction scenario can be dynamically evaluated prior to each control optimization.
Although our validation was conducted on Nordic residential buildings, the adoption of a first-order nodal view for the modeled system, combined with the data-driven, building-specific nature of the covered methods, suggests strong potential for generalizability across building types (e.g., commercial, offices, mixed) and climate zones. Both the reference and proposed model learn parameters directly from building-specific data, allowing them to naturally adapt to different contexts.
@@
The higher computational cost of the LSTM+BNN model is a defined limitation, but we have justified it as manageable through appropriate retraining scheduling and initialization, thus not impeding its practical MPC integration for data-rich buildings. Therefore, integrating the partially stochastic LSTM+BNN into heating MPC solutions—especially for buildings with comprehensive datasets—has great potential to improve control accuracy, thereby enhancing heating control in terms of energy efficiency, thermal comfort, and success with more advanced optimization objectives.
The deployment of an industrial-grade MPC solution in a real-life application is far from trivial and requires significant infrastructure beyond the core model, including data connectivity, storage, robust implementation of control optimization, monitoring, handling connection issues, etc. This work provides a critical building block for such systems by introducing and validating a high-performance, uncertainty-aware model. The integration of this component into a full, robust industrial system remains a separate, important engineering challenge.
@@
Finally, the results of this work motivate and lay the groundwork for further studies on improving probabilistic indoor temperature modeling in the context of heating MPC. Hence, we leave as future work the following:
@@
    Understanding carefully which of the learned physical relationships from the data contribute to the enhanced modeling ability of data-driven models, i.e., identifying what is missing from the existing reference model formulation and how it could be improved. For instance, methods like [55, 56] could be used to find interpretable, functional formulations for the input-output mappings learned by the data-driven models.
@@
    Enhancing the robustness and generalizability of the proposed DL framework. This includes developing more automated and principled methods for hyperparameter selection, such as hierarchical Bayesian modeling, to adapt the prior variances and KL weighting to different building types and data regimes without manual tuning. Furthermore, architectural optimizations for computational efficiency will be crucial for large-scale deployment.
  • (v)
    The logical and necessary next step is the practical implementation and testing within a closed-loop MPC framework. Future work should focus on case studies to evaluate the interaction of the proposed model with different optimization objectives, ensuring control stability. Furthermore, to guarantee robust control actions, the physics-consistency of the model predictions (e.g., monotonicity with respect to certain inputs) must be rigorously studied and enforced. These steps are critical for the large-scale industrial deployment of advanced data-driven models like the one proposed.
@@ 5 Acknowledgments
Appendix A
Algorithm 1 Detailed LSTM+BNN training procedure.
$\bm{X}\in\mathbb{R}^{N\times L\times M},\ \Delta\bm{y}\in\mathbb{R^{N}}$ $\triangleright$ Dataset
$\beta^{2}\leftarrow 10^{-3}$, $w_{\rm KL}\leftarrow 10^{-3}$ $\triangleright$ Set prior variance and KL weight
$p(\bm{z}_{R})\leftarrow\mathcal{N}(\bm{0},\beta^{2}\bm{I})$ $\triangleright$ Define BNN prior
$\bm{\mu}_{R}\leftarrow\bm{0}$, $\bm{\sigma}_{R}\leftarrow\bm{10^{-3}}$, $\bm{z}_{D}\leftarrow\bm{0}$ $\triangleright$ Initialize model parameters
$q\leftarrow\mathcal{N}(\bm{\mu}_{R},\bm{\sigma}_{R}^{2}\bm{I})$ $\triangleright$ Set variational density
$\mathcal{L}_{\text{val,min}}\leftarrow\infty$ $\triangleright$ Initialize minimum validation loss
procedure Training($\bm{X}$)
  $\Delta\bm{y}=\{\Delta\bm{y}_{\text{train}}\in\mathbb{R}^{N_{\text{train}}},\Delta\bm{y}_{\text{val}}\in\mathbb{R}^{N_{\text{val}}}\}$
  for $i=1$ to epochs do
   Draw $\bm{z}_{R}\sim q$ for each $\bm{X}_{\text{train}}^{(t)}$
   With $\Delta\hat{\bm{y}}_{\text{train}},\Delta\bm{y}_{\text{train}},q$ evaluate ELBO as training loss, $\mathcal{L}_{\text{train}}$ $\triangleright$ See equation (16)
   Back-propagate $\mathcal{L}_{\text{train}}$ to compute gradients w.r.t. $\bm{\mu}_{R},\bm{\sigma}_{R},\bm{z}_{D}$
   Update $\bm{\mu}_{R},\bm{\sigma}_{R},\bm{z}_{D}$ based on learning rate and the gradients $\triangleright$ Adam optimizer
   $q\leftarrow\mathcal{N}(\bm{\mu}_{R},\bm{\sigma}^{2}_{R}\bm{I})$ $\triangleright$ Update variational density
   Draw $\bm{z}_{R}\sim q$ for each $\bm{X}_{\text{val}}^{(t)}$
   With $\Delta\hat{\bm{y}}_{\text{val}},\Delta\bm{y}_{\text{val}},q$ evaluate ELBO as validation loss, $\mathcal{L}_{\text{val}}$ $\triangleright$ See equation (16)
   if $\mathcal{L}_{\text{val}}<\mathcal{L}_{\text{val,min}}$ then
     Save model with current parameters $\bm{\mu}_{R},\bm{\sigma}_{R},\bm{z}_{D}$ as the best model, $\Psi_{\text{best}}$
     $\mathcal{L}_{\text{val,min}}\leftarrow\mathcal{L}_{\text{val}}$
  return $\Psi_{\text{best}}$